OpenStudy (anonymous):

expand the expression- log636x2 (the 6 next to the log is the base and the 2 is x squared)

OpenStudy (tkhunny):

\(\log_{6}36^{2} = 2\log_{6}(36) = 2\log_{6}(6^{2}) = 4\log_{6}(6) = 4\) Just rules of logarithms, nothing fancy.

OpenStudy (anonymous):

\[\log_{6}36x^2 = \log_{6}36 + \log_{6}x^2 = 2 + 2\log_{6}x = 2 * (1 + \log_{6}x)\]